# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v2

@@ 2.1 GMV Portfolios under Finite-Sample Risk-Mode Uncertainty
Consider an idealised setting in which an investor observes $\Delta t_{\mathrm{in}}>n$ independent realisations of the $n$-dimensional return vector $\bm{r}_{t}=\left\{r_{t1},r_{t2},\ldots,r_{tn}\right\}^{\top}$. We refer to these observations as the In-Sample (IS) data and assume each $\bm{r}_{t}$ is drawn from a multivariate normal distribution with zero mean and covariance matrix $\bm{\Sigma}$. The task of the GMV portfolio is to find weights $\bm{w}=\left\{w_{1},w_{2},\ldots,w_{n}\right\}^{\top}$ that minimise portfolio variance subject to the full-investment constraint $\bm{1}^{\top}\bm{w}=1$. Replacing the unknown covariance matrix with its maximum-likelihood estimator $\widehat{\bm{\Sigma}}=\bm{\Sigma}_{\text{MLE}}$ yields the typical application of the GMV portfolio
@@
where $q_{\mathrm{in}}=n/\Delta t_{\mathrm{in}}$ and $\sigma_{\star}^{2}=(\bm{1}^{\top}\bm{\Sigma}^{-1}\bm{1})^{-1}$ denotes the population GMV portfolio variance, i.e., an idealized portfolio that is built and tested on the population covariance $\bm{\Sigma}$ [44]. Note that the inflation factor depends only on the estimation window, and it diverges when $\Delta t_{\mathrm{in}}\to n$. Interestingly, $\Delta t_{\mathrm{out}}$ affects only the dispersion of the variance estimator.
@@
In practice, one does not observe the true eigenbasis but only its sample estimate from finite data. RMT results indicate that when a population eigenvalue $\lambda_{k}$ is well separated from the rest of the spectrum, the corresponding sample eigenvector $\widehat{\bm{v}}_{k}$ tends to have a high overlap with the true $\bm{v}_{k}$ [45]. Unfortunately, such spectral isolation occurs primarily for the largest eigenvalues, which enter the global minimum-variance portfolio with weights proportional to $1/\lambda_{k}$ and thus contribute little to risk reduction. The smaller eigenvalues are tightly clustered in the bulk of the spectrum, causing their sample eigenvectors to mix and lose any precise orientation. As a result, estimated eigen-portfolios associated with these directions are dominated by sampling noise rather than true risk modes.
@@
If short selling is forbidden, one adds the element-wise constraint $w_{i}\geq 0$ to the GMV optimization. The problem remains convex but no longer has a closed-form solution; however, it can still be optimally solved by numerical quadratic programming. Empirically, it was observed in Ref. [24, 53] that long-only portfolios on the ML covariance estimator tend to perform similarly to cleaned covariance estimators.
@@ 2.2 Eigenvalue Cleaning for MSE Minimization
$\bm{f}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)=\left\{f_{1}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q),\;\dots,\;f_{n}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)\right\},\quad\text{with}\quad f_{k}:\mathbb{R}^{n+1}\to\mathbb{R};$ (7)
@@
These two structural properties of $\bm{f}$ (Eqs. (7)–(8)) anticipate the constraints we impose on the NN architecture in Sec. 3.2.
The optimal $\bm{f}^{\star}(\widehat{\bm{\lambda}}\mid q)$ is defined as the minimiser of the Frobenius-norm loss against the (unknown) population correlation matrix,
@@
Clearly, this estimator is not available in practice because $\bm{C}$ is unknown. However, under rotational invariance, the expectation of both the population eigenvalues $\bm{\Lambda}$ and the overlap $\bm{\Psi}$ must depend only on the eigenvalues of the empirical covariance matrix $\widehat{\bm{\Lambda}}$ and the aspect ratio $q=n/\Delta t$. Such expectation cannot be derived analytically in a closed form, but NLS theory provides an analytical approximation of $\bm{f}^{\star}\to\bm{f}^{\mathrm{OE}}$ in the limit $n,\Delta t\to\infty$ with $q$ fixed [9].
@@ 3 Neural Networks for Portfolio Optimization
The matrix $\bm{V}_{\mathrm{NN}}$ is obtained from the raw eigenvectors $\widetilde{\bm{V}}$ of the lag-transformed correlation $\widetilde{\bm{C}}$ and projected so that the estimated correlation has ones on its diagonal. In closed form,
@@
3.1 Lag-Transformation Module
@@
with $\bm{\alpha}=\{\alpha_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ and $\bm{\beta}=\{\beta_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ learned by back-propagation. The coefficient $\alpha_{t}$ rescales observations at lag $t$. The coefficient $\beta_{t}$ modulates their outlier sensitivity by controlling the saturation of the hyperbolic tangent. When $\beta_{t}$ is small, the mapping is nearly linear ($\tanh(x)\approx x$ if $x$ is small) and the return enters almost unchanged; when $\beta_{t}$ is large, the mapping approaches a symmetric step that clips extreme values at ±$\pm\alpha_{t}/\beta_{t}$. In this way, the module can emphasise or attenuate individual lags and dampen outliers simultaneously.
The factor 252252252 annualises the input and stabilises gradient computations under single-precision arithmetic. Because $\alpha_{t}$ and $\beta_{t}$ are updated jointly with all other network parameters, the module can learn any monotone or non-monotone lag profile, with the classical exponential decay arising as a special case. The output transformed matrix $\widetilde{\bm{R}}$ serves as the input to all subsequent modules.
@@ 3.2 Eigenvalue Cleaning with Neural Networks
The second module replaces the Ledoit and Péché RIE with a purely data-driven alternative. It learns a vector-valued function $\bm{f}\colon\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ that transforms the empirical eigenvalues $\{\widetilde{\lambda}_{1},\dots,\widetilde{\lambda}_{n}\}$ of the lag-transformed returns’ correlation matrix (Eq. (14)) into the cleaned spectrum $\{\lambda^{\star}_{1},\dots,\lambda^{\star}_{n}\}$.
@@
To clarify the last statement, consider that within any subset of sample eigenvalues, the effective order of interaction is not fixed in advance but grows with the local density of the population spectrum. Regions with tightly spaced population eigenvalues induce strong collective forces that demand modelling higher-order dependencies among neighbouring sample eigenvalues, whereas sparse regions admit accurate approximation with only low-order interactions. In a bottom-up framework such as self-attention, capturing these varying orders requires explicitly stacking as many attention–feed-forward layers as the maximum local interaction order, since each layer only learns pairwise or limited higher-order dependencies. As the dimension $n$ increases, a region of given spectral concentration spans more eigenvalues, forcing deeper networks to maintain the same modelling fidelity. Consequently, this scaling renders pure attention architectures impractical in very high dimensions.
@@
$\bm{g}^{\rightarrow}(\bm{h}^{\rightarrow}_{k-1},\bm{m}^{\rightarrow}_{k-1},\widetilde{\lambda}_{k},q)$ =$=\{\bm{h}^{\rightarrow}_{k},\bm{m}^{\rightarrow}_{k}\},$ $\text{with}\quad k=1,\ldots,n,$ (15)
$\bm{g}^{\leftarrow}(\bm{h}^{\leftarrow}_{k+1},\bm{m}^{\leftarrow}_{k+1},\widetilde{\lambda}_{k},q)$ =$=\{\bm{h}^{\leftarrow}_{k},\bm{m}^{\leftarrow}_{k}\},$ $\text{with}\quad k=n,\ldots,1,$ (16)
@@
In summary, the module comprises a total of $8\omega^{2}+26\omega+1=34,433$ learnable parameters ($\omega=64$). Both LSTM cell contributes $8\omega^{2}+24\omega$ parameters, and the final dense layer adds $2\omega+1$. If model compactness is a priority, the LSTM units may be replaced by two Gated Recurrent Units (GRUs). Standard LSTMs employ separate memory and hidden channels, $\bm{m}$ and $\bm{h}$, respectively, to mitigate vanishing and exploding gradients. The cell state $\bm{m}$, modulated by input, forget, and output gates, can preserve long-range information almost unchanged over many time steps, while the hidden state $\bm{h}$ undergoes more aggressive step-wise updates and serves as the immediate output. This arrangement improves numerical stability and captures both global and local spectral structure. That said, this dual-channel design is not strictly necessary: in our experiments, a simplified GRU, which merges long and short-term memory into a single $\bm{h}$ performed comparably well. The GRU reorganizes its gating mechanism so that the hidden state alone can reliably encode long-range dependencies, offering a lighter yet effective alternative. Substituting LSTM cells with GRUs reduces the total parameter count to $6\omega^{2}+19\omega+1$.
@@ 3.3 Multilayer Perceptron for Inverse Marginal Volatilities
The second branch of the network estimates the inverse marginal volatilities from the sample standard deviation of the transformed return series. i.e.,
@@
Each standard deviation is then passed through an MLP network with 333 hidden layers of (64,32,16)(64,32,16)( 64 , 32 , 16 ) neurons and a Leaky-ReLU activation function. The output of the network is a single neuron with a softplus activation function, which guarantees that the output is positive. In total, this module comprises 2,7532,7532 , 753 learnable parameters. It is important to stress that we process each asset’s estimated volatility independently through the same MLP, taking one scalar at a time rather than feeding the entire vector of $n$ volatilities jointly. This ensures that the network’s parameters remain fixed and do not scale with the number of assets, while still learning a universal transform applicable to every marginal volatility. The output of this branch is then standardized by dividing it by the average inverse output of the network, i.e.,
@@ 4.1 Selection of the Investable Universe
We describe here the procedure used to select assets for investment from the universe of US equities that were listed on the NYSE or NASDAQ at any point from 1 January 1990 to 31 December 2024. This universe, therefore, includes American Depositary Receipts (ADRs) and explicitly excludes funds and exchange-traded funds. To avoid data leakage, for each trading day $t$, we rely only on information available up to $t-1$ when determining the eligible stocks. We further exclude any security for which a delisting occurs within the subsequent $\Delta t_{\text{out}}=5$ trading days. This procedure is safe because, under SEC rules, issuers must provide at least 10 days’ written notice to the exchange before filing Form 25 to delist securities [63].
We designed a two-tiered filtering scheme that operates on both the full five-year calibration window ($\Delta t_{\text{in}}=1200$) and a short, recent time window. The full-history filters enforce homogeneity of trading activity over the entire five years and thus attempt to remove any stock whose price series is fragmented by late initial public offerings, prolonged trading halts, or significant corporate restructurings. The recent-window filters, by contrast, apply stringent liquidity thresholds over the last five to twenty trading days with the attempt to eliminate any security whose executed auctions fall below the levels compatible with low-impact rebalancing. This dual approach both addresses the challenges of assembling a broad universe of $n=1000$ stocks with uninterrupted five-year histories and ensures that each selected asset remains sufficiently liquid at the time of investment. That said, the threshold values themselves are inherently arbitrary and have been chosen primarily to maintain a stable universe of $n=1000$ stocks at every point in time; nevertheless, filtering approaches like this reflect standard practice in hedge funds, where inclusion criteria must satisfy stringent and complex rules. In practice, working with a very large, unconstrained dataset can introduce unintended biases and degrade the reliability of both risk estimates and optimized allocations.
@@
Finally, to reduce redundancy, we keep only one class of shares for each company, the one with the largest market capitalization on the previous day. A similar filter is executed for those stocks with an in-sample correlation greater than 0.95 [65]. The final selection is done by ranking the stocks by their market capitalization and selecting the top $n=1,000$ stocks for every trading day $t$. The final panel, spanning January 1990 to December 2024, contains roughly 2.72.72.7 millions of stock–day pairs and approximately 4,0004,0004 , 000 unique stocks.

### v3

@@ 2.1 GMV Portfolios under Finite-Sample Risk-Mode Uncertainty
The section motivates why finite-sample estimation error inflates OOS portfolio risk in high dimensions and interprets this effect in terms of noisy risk modes. For analytical tractability, we show a closed-form reference expression for the resulting risk inflation under multivariate normal and i.i.d. returns. These assumptions are used only for the benchmark formula and are not required by our covariance-filtering method or by the empirical analysis, which is conducted on real equity returns that are non-Gaussian and exhibit heavy tails and conditional heteroskedasticity.
Consider an idealised setting in which an investor observes $\Delta t_{\mathrm{in}}>n$ independent realisations of the $n$-dimensional return vector $\bm{r}_{t}=\left\{r_{t1},r_{t2},\ldots,r_{tn}\right\}^{\top}$. We refer to these observations $\Delta t_{\mathrm{in}}$ as the In-Sample (IS) data and assume each $\bm{r}_{t}$ is drawn from a multivariate normal distribution with zero mean and covariance matrix $\bm{\Sigma}$. The task of the GMV portfolio is to find weights $\bm{w}=\left\{w_{1},w_{2},\ldots,w_{n}\right\}^{\top}$ that minimise portfolio variance subject to the full-investment constraint $\bm{1}^{\top}\bm{w}=1$. Replacing the unknown covariance matrix with its Maximum-Likelihood Estimator (MLE) $\widehat{\bm{\Sigma}}=\bm{\Sigma}_{\text{MLE}}$ yields the typical application of the GMV portfolio
@@
where $q_{\mathrm{in}}:=n/\Delta t_{\mathrm{in}}$ and $\sigma_{\star}^{2}=(\bm{1}^{\top}\bm{\Sigma}^{-1}\bm{1})^{-1}$ denotes the population GMV portfolio variance, i.e., an idealized portfolio that is built and tested on the population covariance $\bm{\Sigma}$ [29]. Note that the inflation factor depends only on the estimation window, and it diverges when $\Delta t_{\mathrm{in}}\to n$. Interestingly, $\Delta t_{\mathrm{out}}$ affects only the dispersion of the variance estimator.
Eq. (2) highlights that, even under an idealised stationary Gaussian setting, OOS risk inflation is governed by the aspect ratio $q_{\mathrm{in}}=n/\Delta t_{\mathrm{in}}$. In a stationary environment, the standard remedy is to reduce $q_{\mathrm{in}}$ by increasing the estimation window $\Delta t_{\mathrm{in}}$, thereby letting sampling noise average out. For financial covariances, however, this prescription is intrinsically limited because the dependence structure evolves over time: enlarging $\Delta t_{\mathrm{in}}$ mixes heterogeneous market regimes and yields an estimator that is increasingly biased with respect to the covariance relevant for the subsequent holding period. In other words, non-stationarity limit the effective look-back horizon, because observations that are too old become statistically outdated for forecasting near-term risk.
This creates a structural tension in large-scale portfolio construction. Increasing $n$ is attractive because it enlarges the diversification set and may lower the population GMV variance $\sigma_{\star}^{2}$, yet non-stationarity constrains $\Delta t_{\mathrm{in}}$ to remain short, so $q_{\mathrm{in}}:=n/\Delta t_{\mathrm{in}}$ grows and Eq. (2) predicts stronger finite-sample risk inflation. In this regime, the usual stabilisation strategy of extending $\Delta t_{\mathrm{in}}$ is problematic: it reduces sampling variance but increases bias by mixing heterogeneous regimes, whereas shortening $\Delta t_{\mathrm{in}}$ limits this bias but makes $\widehat{\bm{\Sigma}}$ ill-conditioned and $\widehat{\bm{\Sigma}}^{-1}$ unstable. Our regularisation should therefore be read as a pragmatic response to this constraint: it does not forecast regime shifts, but rather emphasises the most recent information and aims to extract a more reliable covariance signal from limited, potentially more relevant samples.
@@
In practice, one does not observe the true eigenbasis but only its sample estimate from finite data. RMT results indicate that when a population eigenvalue $\lambda_{k}$ is well separated from the rest of the spectrum, the corresponding sample eigenvector $\widehat{\bm{v}}_{k}$ tends to have a high overlap with the true $\bm{v}_{k}$ [62]. Unfortunately, such spectral isolation occurs primarily for the largest eigenvalues, which enter the global minimum‐variance portfolio with weights proportional to $1/\lambda_{k}$ and thus contribute little to risk reduction. The smaller eigenvalues are tightly clustered in the bulk of the spectrum, causing their sample eigenvectors to mix and lose any precise orientation. As a result, estimated eigen‐portfolios associated with these directions are dominated by sampling noise rather than true risk modes.
@@
If short selling is forbidden, one adds the element-wise constraint $w_{i}\geq 0$ to the GMV optimization. The problem remains convex but no longer has a closed-form solution; however, it can still be optimally solved by numerical QP. Empirically, prior studies have reported that, under long-only constraints, portfolios constructed using the sample covariance matrix (i.e., the MLE) often deliver performance comparable to that obtained with covariance cleaning procedures; see Refs. [63, 39].
@@ 2.2 Eigenvalue Cleaning for MSE Minimization
Because the map depends only on the empirical spectrum (and $q$), it is independent of the empirical eigenvectors and therefore basis-independent: any orthogonal similarity transform $\mathbf{O}\bm{C}_{\mathrm{MLE}}\mathbf{O}^{\top}$ yields the same eigenvalue correction, with the cleaned matrix rotating accordingly.
@@
$\bm{f}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)=\left\{f_{1}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q),\;\dots,\;f_{n}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)\right\},\\ \text{with}\quad f_{k}:\mathbb{R}^{n+1}\to\mathbb{R};$ (7)
@@
Here $\tau$ denotes a relabelling of eigenpairs. Permuting the entries of $\widehat{\bm{\lambda}}$ is implicitly accompanied by the same permutation of the associated eigenvectors (i.e., a column permutation of $\widehat{\mathbf{V}}$), so that the eigendecomposition of $\widehat{\mathbf{C}}$ is left unchanged. We note that Eq. (8) does not introduce any additional probabilistic assumption beyond the standard RIE framework. The sample eigendecomposition is defined only up to a relabelling of eigenpairs; therefore, any estimator that acts on the empirical spectrum and keeps the empirical eigenvectors must be consistent under such relabellings, which is exactly the permutation-quivariance property stated in Eq. (8). These two structural properties of $\bm{f}$ (Eqs. (7)–(8)) anticipate the constraints we impose on the NN architecture in Sec. 3.2.
The optimal $\bm{f}^{\star}(\widehat{\bm{\lambda}}\mid q)$ is defined as the minimiser of the Frobenius‐norm loss against the (unknown) population correlation matrix,
@@
Clearly, this estimator is not available in practice because $\bm{C}$ is unknown. However, under rotational invariance, the expectation of both the population eigenvalues $\bm{\Lambda}$ and the overlap $\bm{\Psi}$ must depend only on the eigenvalues of the empirical covariance matrix $\widehat{\bm{\Lambda}}$ and the aspect ratio $q:=n/\Delta t_{\textrm{in}}$. Such expectation cannot be derived analytically in a closed form, but NLS theory provides an analytical approximation of $\bm{f}^{\star}\to\bm{f}^{\mathrm{OE}}$ in the limit $n,\Delta t_{\textrm{in}}\to\infty$ with $q$ fixed [49].
@@ 3 Neural Networks for Portfolio Optimization
The matrix $\bm{V}_{\mathrm{NN}}$ is obtained from the raw eigenvectors $\widetilde{\bm{V}}$ of the lag‐transformed correlation $\widetilde{\bm{C}}$ and projected so that the estimated correlation has ones on its diagonal. In closed form,
@@
A reference implementation is available online and can be installed via PyPI [12].
3.1 Lag‐Transformation Module
@@
with $\bm{\alpha}=\{\alpha_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ and $\bm{\beta}=\{\beta_{t}\}_{t=1}^{\Delta t_{\text{in}}}$ learned end-to-end together with the rest of the pipeline by back-propagation. The specific functional form in Eq. (14) is primarily motivated by monotonicity and smooth differentiability, which supports stable back-propagation, by boundedness, which yields robustness to outliers through saturation, and by interpretability, since $\alpha_{t}$ controls the local linear scale while $\alpha_{t}/\beta_{t}$ sets the saturation level. This choice is not unique, and alternative monotone, differentiable, bounded parametrizations could be used without changing the overall framework.. Moreover, we stress that Eq. (14) defines an element-wise differentiable parametric transform, not a feed-forward Multi-Layer Perceptron (MLP). The coefficient $\alpha_{t}$ rescales observations at lag $t$. The coefficient $\beta_{t}$ modulates their outlier sensitivity by controlling the saturation of the hyperbolic tangent. When $\beta_{t}$ is small, the mapping is nearly linear ($\tanh(x)\approx x$ if $x$ is small) and the return enters almost unchanged; when $\beta_{t}$ is large, the mapping approaches a symmetric step that clips extreme values at $\pm\alpha_{t}/\beta_{t}$. In this way, the module can emphasise or attenuate individual lags and dampen outliers simultaneously.
The factor $252$ annualises the input and stabilises gradient computations under single-precision arithmetic. Because $\alpha_{t}$ and $\beta_{t}$ are updated jointly with all other network parameters, the module can learn any monotone or non-monotone lag profile, with the classical exponential decay arising as a special case. The output transformed matrix $\widetilde{\bm{R}}$ serves as the input to all subsequent modules.
@@ 3.2 Eigenvalue Cleaning with Neural Networks
The second module replaces the Ledoit and Péché RIE with a purely data-driven alternative. It learns a vector-valued function $\bm{f}\colon\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ that transforms the empirical eigenvalues $\{\widetilde{\lambda}_{1},\dots,\widetilde{\lambda}_{n}\}$ of the lag-transformed returns’ correlation matrix (Eq. (14)) into the cleaned spectrum $\{\lambda^{\star}_{1},\dots,\lambda^{\star}_{n}\}$. Since the learned filtering depends only on the empirical eigenvalues, it is independent of the associated eigenvectors and hence basis-independent: any orthogonal similarity transform of the input correlation matrix yields the same eigenvalue correction, with the cleaned matrix rotating accordingly.
@@
To clarify the last statement, consider that within any subset of sample eigenvalues, the effective order of interaction is not fixed in advance but grows with the local density of the population spectrum. Regions with tightly spaced population eigenvalues induce strong collective forces that demand modelling higher‐order dependencies among neighbouring sample eigenvalues, whereas sparse regions admit accurate approximation with only low‐order interactions. In a bottom‐up framework such as self‐attention, capturing these varying orders requires explicitly stacking as many attention–feed‐forward layers as the maximum local interaction order, since each layer only learns pairwise or limited higher‐order dependencies. As the dimension $n$ increases, a region of given spectral concentration spans more eigenvalues, forcing deeper networks to maintain the same modelling fidelity. Consequently, this scaling renders pure attention architectures impractical in very high dimensions.
@@
$\bm{g}^{\rightarrow}(\bm{h}^{\rightarrow}_{k-1},\bm{m}^{\rightarrow}_{k-1},\widetilde{\lambda}_{k},q)$ $=\{\bm{h}^{\rightarrow}_{k},\bm{m}^{\rightarrow}_{k}\},$ $\text{with}\quad k=1,\ldots,n,$ (15)
$\bm{g}^{\leftarrow}(\bm{h}^{\leftarrow}_{k+1},\bm{m}^{\leftarrow}_{k+1},\widetilde{\lambda}_{k},q)$ $=\{\bm{h}^{\leftarrow}_{k},\bm{m}^{\leftarrow}_{k}\},$ $\text{with}\quad k=n,\ldots,1,$ (16)
@@
In summary, the module comprises a total of $8\omega^{2}+26\omega+1=34,433$ learnable parameters ($\omega=64$). Both LSTM cell contributes $8\omega^{2}+24\omega$ parameters, and the final dense layer adds $2\omega+1$. If model compactness is a priority, the LSTM units may be replaced by two Gated Recurrent Units (GRUs). Standard LSTMs employ separate memory and hidden channels, $\bm{m}$ and $\bm{h}$, respectively, to mitigate vanishing and exploding gradients. The cell state $\bm{m}$, modulated by input, forget, and output gates, can preserve long-range information almost unchanged over many time steps, while the hidden state $\bm{h}$ undergoes more aggressive step-wise updates and serves as the immediate output. This arrangement improves numerical stability and captures both global and local spectral structure. That said, this dual‐channel design is not strictly necessary: in our experiments, a simplified GRU, which merges long and short‐term memory into a single $\bm{h}$ performed comparably well. The GRU reorganizes its gating mechanism so that the hidden state alone can reliably encode long-range dependencies, offering a lighter yet effective alternative. Substituting LSTM cells with GRUs reduces the total parameter count to $6\omega^{2}+19\omega+1$.
@@ 3.3 Multilayer Perceptron for Inverse Marginal Volatilities
The second branch of the network estimates the inverse marginal volatilities from the sample standard deviation of the transformed return series (Eq. (14)). i.e.,
@@
Each standard deviation is then passed through an MLP network with $3$ hidden layers of $(64,32,16)$ neurons and a Leaky-ReLU activation function. The output of the network is a single neuron with a softplus activation function, which guarantees that the output is positive. In total, this module comprises $2,753$ learnable parameters. It is important to stress that we process each asset’s estimated volatility independently through the same MLP, taking one scalar at a time rather than feeding the entire vector of $n$ volatilities jointly. This ensures that the network’s parameters remain fixed and do not scale with the number of assets, while still learning a universal transform applicable to every marginal volatility. The output of this branch is then standardized by dividing it by the average inverse output of the network, i.e.,
@@ 4.1 Selection of the Investable Universe
We describe here the procedure used to select assets for investment from the universe of US equities that were listed on the NYSE or NASDAQ at any point from 1 January 1990 to 31 December 2024. This universe, therefore, includes American Depositary Receipts (ADRs) and explicitly excludes funds and exchange-traded funds. To avoid data leakage, for each trading day $t$, we rely only on information available up to t1t-1 when determining the eligible stocks. We further exclude any security for which a delisting occurs within the subsequent $\Delta t_{\text{out}}=5$ trading days. This procedure is safe because, under SEC rules, issuers must provide at least 10 days’ written notice to the exchange before filing Form 25 to delist securities [74].
We designed a two‐tiered filtering scheme that operates on both the full five-year calibration window ($\Delta t_{\text{in}}=1200$) and a short, recent time window. The full-history filters enforce homogeneity of trading activity over the entire five years and thus attempt to remove any stock whose price series is fragmented by late initial public offerings, prolonged trading halts, or significant corporate restructurings. The recent-window filters, by contrast, apply stringent liquidity thresholds over the last five to twenty trading days with the attempt to eliminate any security whose executed auctions fall below the levels compatible with low-impact rebalancing. This dual approach both addresses the challenges of assembling a broad universe of $n=1000$ stocks with uninterrupted five-year histories and ensures that each selected asset remains sufficiently liquid at the time of investment. That said, the threshold values themselves are inherently arbitrary and have been chosen primarily to maintain a stable universe of $n=1000$ stocks at every point in time; nevertheless, filtering approaches like this reflect standard practice in hedge funds, where inclusion criteria must satisfy stringent and complex rules. In practice, working with a very large, unconstrained dataset can introduce unintended biases and degrade the reliability of both risk estimates and optimized allocations.
@@
Finally, to reduce redundancy, we keep only one class of shares for each company, the one with the largest market capitalization on the previous day. A similar filter is executed for those stocks with an in-sample correlation greater than 0.95 [21]. The final selection is done by ranking the stocks by their market capitalization and selecting the top $n=1,000$ stocks for every trading day $t$. The final panel, spanning January 1990 to December 2024, contains roughly $2.7$ millions of stock–day pairs and approximately $4,000$ unique stocks.